# 3D and 4D World Modeling: A Survey

### v2

@@ 4.2.4 Reconstruction-Centric Quality
4.2.5 Downstream Evaluation
4.3 Quantitative Experiments & Analyses
4.3.1 Benchmarking Video Generation Models
Generation Fidelity. Table VI reports FID and FVD results on the nuScenes validation set for both single-view and multi-view vision-based world models. Early baselines such as GenAD [96] and DriveDreamer [58] operate at relatively low resolutions and frame rates, achieving modest performance (FID $\sim$15, FVD 180–340). Later single-view models improve visual quality. Vista [103] and InfinityDrive [107] leverage higher resolutions and frame rates, reducing FVD below 100. Recent works like MaskGWM [117] and GeoDrive [94] set new state-of-the-art, reaching FID around 4–5 and FVD near 60. In the multi-view setting, early BEV-based approaches (BEVControl [73], BEVGen [72]) yield high FID (>>20). Subsequent models, including DriveWM [102], Panacea [80], and MagicDrive [20] reduce errors but struggle with temporal stability (FVD >>120). Strong improvements come from models emphasizing geometric consistency and spatio-temporal alignment. UniScene [77], DriveScape [114], and DiST-4D [121] achieve the best balance, with FVD scores below 80 and DiST-4D [121] reaching as low as 22.67.
Downstream Evaluations. Table VIII and Table VII evaluate downstream perception and planning on generated scenes. Early generative baselines (BEVControl [73], BEVGen [72]) provide limited perception benefits, especially in vehicle segmentation ($27\%$ mIoU). More advanced methods such as MagicDrive [20] and DreamForge [105] improve both detection (up to 26 mAP on StreamPETR [256]) and segmentation ($61\%$ road mIoU), while DrivePhysica [93] and Glad [83] further push detection accuracy (35.5 mAP, 43.7 NDS). For segmentation, UniMLVG [92] and CogDriving [108] achieve the highest fidelity (70.8% road, 32.1% vehicle mIoU). Beyond perception, planning performance highlights the persistent gap between synthetic and real data. While real nuScenes provides the upper bound (37.9 mAP, 49.9 NDS, 1.05 Avg L2), generative methods lag significantly in detection and planning accuracy. Nevertheless, world models like DriveArena [119] and DreamForge demonstrate reduced planning errors and collision rates, enabling preliminary closed-loop driving with non-trivial success rates (e.g., 0.81 PDMS). DrivingSphere [115] achieves the strongest drivable area segmentation (>>58% mIoU), while DiST-4D [121] balances detection and segmentation performance but lacks closed-loop validation.
4.3.2 Benchmarking Occupancy Generation Models
Occupancy Reconstruction Quality. Table IX evaluates the reconstruction capability of occupancy world models under VAE-based formulations. Conventional VAEs such as DOME [183] already achieve strong results (83.08% mIoU, 77.25% IoU), outperforming most VQVAEs. While UrbanDiff [178] and I2World [195] show competitive IoU, other variants like OccSora [179] degrade significantly under coarse temporal–spatial compression. Triplane-based VAEs [173, 196, 198] bring the largest gains, with T3Former [196] reaching 85.50% mIoU and X-Scene [198] establishing a new state-of-the-art at 92.40% mIoU and 85.60% IoU.
TABLE IX: Benchmarking OccGen models on Reconstruction Quality. The reported metrics are mIoU (%) for Semantic Occupancy Reconstruction and IoU (%) for Occupancy Reconstruction. All results are on the official nuScenes [10] validation set. Both metrics are the higher the better ($\uparrow$).
@@ 4.3.2 Benchmarking Occupancy Generation Models
TABLE X: Benchmarking OccGen models on 4D Occupancy Forecasting Quality. The reported metrics are mIoU (%) for Semantic Occupancy Reconstruction and IoU (%) for Occupancy Reconstruction, respectively, at timestamps 1s, 2s, and 3s. All results are on the official nuScenes [10] validation set. Both metrics are the higher the better ($\uparrow$).
@@
TABLE XI: Benchmarking OccGen models on Motion Planning Quality. The reported metrics are L2 Error Rate (in meters) and Collision Rate (%), respectively, at timestamps 1s, 2s, and 3s. All results are on the official nuScenes [10] validation set. Both metrics are the lower the better ($\downarrow$).
@@
4D Occupancy Forecasting Quality. Table X presents 4D occupancy forecasting results over the period of 1–3 seconds. Baselines such as OccWorld [177] and OccLLaMA [181] achieve moderate performance (17–20% mIoU), while DOME [183] and UniScene [77] improve temporal stability (27.10% and 31.76% mIoU). More recent models show further progress: I2World [195] reaches 39.73% mIoU with balanced IoU, and T3Former [196] excels in spatial coherence with 76.40% IoU. The comparisons reveal three insights. First, naive autoregressive or generative approaches deteriorate rapidly at longer horizons, highlighting the need for structured priors. Second, triplane factorization substantially improves spatial fidelity, as reflected in the performance of T3Former [196]. Third, I2World shows that coupling scalable latent reasoning with temporal modeling yields the best balance across horizons. Accurate 4D forecasting thus requires not only generative power but also structured representations that enforce geometric and temporal consistency.
End-to-End Planning. Table XI reports the performance of end-to-end planning, measured by trajectory error (L2) and collision rate. Sequence-based planners like ST-P3 [263] perform poorly (2.11 meters in L2 error), while UniAD [254] and GenAD [96] achieve substantial gains, with UniAD+DriveWorld [174] further improving to 0.69 meter in L2 error and 0.19% collisions. Occupancy-based world models such as OccWorld [177] and OccLLaMA [181] reduce errors to around 1.15 meters. Structured refinements (e.g., DFIT-OccWorld [185], RenderWorld [189], and Drive-OccWorld [186]) achieve stronger accuracy and safety, with Drive-OccWorld reaching 0.85 m in L2 error and 0.29% collisions. Notably, GaussianAD [184] and T3Former [196] balance error and safety, while Occ-LLM [190] reports extremely low error (i.e., only 0.28 meter in L2 error). The results show that integrating occupancy world models into planning pipelines consistently outperforms pure trajectory-based methods. Hybrid designs that refine occupancy priors, such as Drive-OccWorld [186] and DFIT-OccWorld [185], bring joint improvements in accuracy and safety, demonstrating the downstream robustness of generative modeling. Overall, structured occupancy representations form a strong foundation for end-to-end autonomous driving, enabling reliable long-horizon planning in complex scenarios.
4.3.3 Benchmarking LiDAR Generation Models
Generation Fidelity. Table XII reports the performance of recent LiDAR scene generation methods on SemanticKITTI [16] using four fidelity metrics (FRD, FPD, JSD, and MMD). Earlier methods such as LiDARGen [205] and LiDM [209] exhibit relatively large distributional discrepancies, as reflected by high FRD and FPD scores. In contrast, more recent approaches, including R2DM [206], Text2LiDAR [223], and WeatherGen [229], achieve substantially better results across most metrics, indicating a closer alignment between generated and real LiDAR distributions.
TABLE XII: Benchmarking LiDARGen models on the Perceptual Fidelity evaluations. The reported metrics are FRD, FPD, JSD and MMD scores on the official SemanticKITTI [16] dataset. All metrics are the lower the better ($\downarrow$).
@@ 4.3.3 Benchmarking LiDAR Generation Models
TABLE XIII: Benchmarking LiDARGen models on 4D LiDAR Generation Quality. The reported metrics are TTCE and CTC. The numbers indicate frame intervals. All results are on nuScenes [10]. Both metrics are the lower the better ($\downarrow$).
@@
4D LiDAR Generation Quality. Table XIII benchmarks recent LiDAR-based 4D scene generation methods on temporal coherence, using TTCE (Temporal Transformation Consistency Error) and CTC (Chamfer Temporal Consistency) as evaluation metrics. Unlike video generation, which has been extensively studied with standardized benchmarks, temporal LiDAR generation remains relatively underexplored, and current metrics mainly focus on explicit geometric alignment across frames. The results reveal several observations. First, end-to-end autoregressive methods such as UniScene [77] and OpenDWM-DiT [236] demonstrate clear advantages in maintaining short-horizon geometric consistency, as reflected in lower TTCE and CTC at 1–2 frame intervals. However, their fixed-length generation limits broader applicability, as error accumulation grows at longer horizons. Second, incorporating strong vector quantization modules [236] facilitates better condition embedding and fine-grained reconstruction, leading to improved temporal stability. Third, modality choices introduce inherent trade-offs: BEV-based generation offers smoother temporal continuity but sacrifices fidelity to the raw point cloud pattern, while range-based [49] generation better preserves LiDAR-specific sensing characteristics but requires careful design to embed conditions and sustain long-term consistency.
@@ 4.4 Qualitative Experiments & Analyses
Figure 9: Qualitative examples of OccGen models on the nuScenes [10] dataset. From left to right columns: The input condition, the generated multi-view videos, and the generated occupancy grids. The results are generated using $\mathcal{X}$-Scene [198].
@@ 4.4.2 Qualitative Analyses of OccGen Models
3D Geometric Consistency. Figure 9 shows qualitative results of occupancy generation models conditioned on scene layouts. The generated multi-view videos and occupancies exhibit strong spatial alignment across different perspectives. Such cross-view coherence is crucial for maintaining geometric plausibility in multi-camera settings.
@@ 7 Conclusion
TABLE XIV: Summary of the evaluation metrics used for evaluating the quality of 1Generation, 2Forecasting, 3Planning, 4Reconstruction, and 5Downstream Tasks for the VideoGen, OccGen, and LiDARGen models in 2D, 3D, and 4D tasks.

### v3

@@ 4.2.4 Reconstruction-Centric Quality
Table 9: Benchmarking OccGen models on Reconstruction Quality. The reported metrics are mIoU (%) for Semantic Occupancy Reconstruction and IoU (%) for Occupancy Reconstruction. All results are on the official nuScenes [24] validation set. Both metrics are the higher the better ($\uparrow$).
@@
4.2.5 Downstream Evaluation
Table 10: Benchmarking OccGen models on 4D Occupancy Forecasting Quality. The reported metrics are mIoU (%) for Semantic Occupancy Reconstruction and IoU (%) for Occupancy Reconstruction, respectively, at timestamps 1s, 2s, and 3s. All results are on the official nuScenes [24] validation set. Both metrics are the higher the better ($\uparrow$).
@@ 4.2.5 Downstream Evaluation
4.3 Quantitative Experiments & Analyses
4.3.1 Benchmarking Video Generation Models
Generation Fidelity. Table 6 reports FID and FVD results on the nuScenes validation set for both single-view and multi-view vision-based world models. Early baselines such as GenAD [274] and DriveDreamer [236] operate at relatively low resolutions and frame rates, achieving modest performance (FID $\sim$15, FVD 180–340). Later single-view models improve visual quality. Vista [65] and InfinityDrive [77] leverage higher resolutions and frame rates, reducing FVD below 100. Recent works like MaskGWM [178] and GeoDrive [31] set new state-of-the-art, reaching FID around 4–5 and FVD near 60. In the multi-view setting, early BEV-based approaches (BEVControl [275], BEVGen [211]) yield high FID (>>20). Subsequent models, including DriveWM [239], Panacea [244], and MagicDrive [64] reduce errors but struggle with temporal stability (FVD >>120). Strong improvements come from models emphasizing geometric consistency and spatio-temporal alignment. UniScene [125], DriveScape [249], and DiST-4D [76] achieve the best balance, with FVD scores below 80 and DiST-4D [76] reaching as low as 22.67.
Table 11: Benchmarking OccGen models on Motion Planning Quality. The reported metrics are L2 Error Rate (in meters) and Collision Rate (%), respectively, at timestamps 1s, 2s, and 3s. All results are on the official nuScenes [24] validation set. Both metrics are the lower the better ($\downarrow$).
@@ 4.3.1 Benchmarking Video Generation Models
Downstream Evaluations. Table 8 and Table 7 evaluate downstream perception and planning on generated scenes. Early generative baselines (BEVControl [275], BEVGen [211]) provide limited perception benefits, especially in vehicle segmentation ($27\%$ mIoU). More advanced methods such as MagicDrive [64] and DreamForge [163] improve both detection (up to 26 mAP on StreamPETR [231]) and segmentation ($61\%$ road mIoU), while DrivePhysica [286] and Glad [256] further push detection accuracy (35.5 mAP, 43.7 NDS). For segmentation, UniMLVG [34] and CogDriving [153] achieve the highest fidelity (70.8% road, 32.1% vehicle mIoU). Beyond perception, planning performance highlights the persistent gap between synthetic and real data. While real nuScenes provides the upper bound (37.9 mAP, 49.9 NDS, 1.05 Avg L2), generative methods lag significantly in detection and planning accuracy. Nevertheless, world models like DriveArena [278] and DreamForge demonstrate reduced planning errors and collision rates, enabling preliminary closed-loop driving with non-trivial success rates (e.g., 0.81 PDMS). DrivingSphere [268] achieves the strongest drivable area segmentation (>>58% mIoU), while DiST-4D [76] balances detection and segmentation performance but lacks closed-loop validation.
4.3.2 Benchmarking Occupancy Generation Models
Occupancy Reconstruction Quality. Table 9 evaluates the reconstruction capability of occupancy world models under VAE-based formulations. Conventional VAEs such as DOME [72] already achieve strong results (83.08% mIoU, 77.25% IoU), outperforming most VQVAEs. While UrbanDiff [294] and I2World [142] show competitive IoU, other variants like OccSora [229] degrade significantly under coarse temporal–spatial compression. Triplane-based VAEs [124, 261, 280] bring the largest gains, with T3Former [261] reaching 85.50% mIoU and X-Scene [280] establishing a new state-of-the-art at 92.40% mIoU and 85.60% IoU.
Table 12: Benchmarking LiDARGen models on the Perceptual Fidelity evaluations. The reported metrics are FRD, FPD, JSD and MMD scores on the official SemanticKITTI [14] dataset. All metrics are the lower the better ($\downarrow$).
@@ 4.3.2 Benchmarking Occupancy Generation Models
4D Occupancy Forecasting Quality. Table 10 presents 4D occupancy forecasting results over the period of 1–3 seconds. Baselines such as OccWorld [308] and OccLLaMA [242] achieve moderate performance (17–20% mIoU), while DOME [72] and UniScene [125] improve temporal stability (27.10% and 31.76% mIoU). More recent models show further progress: I2World [142] reaches 39.73% mIoU with balanced IoU, and T3Former [261] excels in spatial coherence with 76.40% IoU. The comparisons reveal three insights. First, naive autoregressive or generative approaches deteriorate rapidly at longer horizons, highlighting the need for structured priors. Second, triplane factorization substantially improves spatial fidelity, as reflected in the performance of T3Former [261]. Third, I2World shows that coupling scalable latent reasoning with temporal modeling yields the best balance across horizons. Accurate 4D forecasting thus requires not only generative power but also structured representations that enforce geometric and temporal consistency.
Table 13: Benchmarking LiDARGen models on 4D LiDAR Generation Quality. The reported metrics are TTCE and CTC. The numbers indicate frame intervals. All results are on nuScenes [24]. Both metrics are the lower the better ($\downarrow$).
@@
End-to-End Planning. Table 11 reports the performance of end-to-end planning, measured by trajectory error (L2) and collision rate. Sequence-based planners like ST-P3 [90] perform poorly (2.11 meters in L2 error), while UniAD [92] and GenAD [274] achieve substantial gains, with UniAD+DriveWorld [169] further improving to 0.69 meter in L2 error and 0.19% collisions. Occupancy-based world models such as OccWorld [308] and OccLLaMA [242] reduce errors to around 1.15 meters. Structured refinements (e.g., DFIT-OccWorld [292], RenderWorld [273], and Drive-OccWorld [281]) achieve stronger accuracy and safety, with Drive-OccWorld reaching 0.85 m in L2 error and 0.29% collisions. Notably, GaussianAD [310] and T3Former [261] balance error and safety, while Occ-LLM [265] reports extremely low error (i.e., only 0.28 meter in L2 error). The results show that integrating occupancy world models into planning pipelines consistently outperforms pure trajectory-based methods. Hybrid designs that refine occupancy priors, such as Drive-OccWorld [281] and DFIT-OccWorld [292], bring joint improvements in accuracy and safety, demonstrating the downstream robustness of generative modeling. Overall, structured occupancy representations form a strong foundation for end-to-end autonomous driving, enabling reliable long-horizon planning in complex scenarios.
4.3.3 Benchmarking LiDAR Generation Models
Generation Fidelity. Table 12 reports the performance of recent LiDAR scene generation methods on SemanticKITTI [14] using four fidelity metrics (FRD, FPD, JSD, and MMD). Earlier methods such as LiDARGen [324] and LiDM [192] exhibit relatively large distributional discrepancies, as reflected by high FRD and FPD scores. In contrast, more recent approaches, including R2DM [173], Text2LiDAR [250], and WeatherGen [251], achieve substantially better results across most metrics, indicating a closer alignment between generated and real LiDAR distributions.
4D LiDAR Generation Quality. Table 13 benchmarks recent LiDAR-based 4D scene generation methods on temporal coherence, using TTCE (Temporal Transformation Consistency Error) and CTC (Chamfer Temporal Consistency) as evaluation metrics. Unlike video generation, which has been extensively studied with standardized benchmarks, temporal LiDAR generation remains relatively underexplored, and current metrics mainly focus on explicit geometric alignment across frames. The results reveal several observations. First, end-to-end autoregressive methods such as UniScene [125] and OpenDWM-DiT [200] demonstrate clear advantages in maintaining short-horizon geometric consistency, as reflected in lower TTCE and CTC at 1–2 frame intervals. However, their fixed-length generation limits broader applicability, as error accumulation grows at longer horizons. Second, incorporating strong vector quantization modules [200] facilitates better condition embedding and fine-grained reconstruction, leading to improved temporal stability. Third, modality choices introduce inherent trade-offs: BEV-based generation offers smoother temporal continuity but sacrifices fidelity to the raw point cloud pattern, while range-based [139] generation better preserves LiDAR-specific sensing characteristics but requires careful design to embed conditions and sustain long-term consistency.
@@ 4.4 Qualitative Experiments & Analyses
Figure 9: Qualitative examples of OccGen models on nuScenes [24]. From left to right columns: The input condition, the generated multi-view videos, and the generated occupancy grids. The results are generated using $\mathcal{X}$-Scene [280].
@@ 4.4.2 Qualitative Analyses of OccGen Models
3D Geometric Consistency. Figure 9 shows qualitative results of occupancy generation models conditioned on scene layouts. The generated videos and occupancies exhibit strong spatial alignment across perspectives. Such cross-view coherence is crucial for maintaining geometric plausibility in multi-camera settings.
@@ 7 Conclusion
Table 14: Summary of the evaluation metrics used for evaluating the quality of 1Generation, 2Forecasting, 3Planning, 4Reconstruction, and 5Downstream Tasks for the VideoGen, OccGen, and LiDARGen models in 2D, 3D, and 4D tasks.